There are three forces 4n 6n and 8n can their resultant be equal to zero

Grade:11

1 Answers

Sarfaraz Ahamed
6 years ago
Yes, it is possible. One of the several ways possible is through the way mentioned below:
when 4N and 6N forces are kept at some angle where by their resultant is 8N force.
If this 8N force is in opposite dirction to the 8N force (given) it can be a zero vector.
 
Let us find the angle between 4N and 6N forces so that we can get a 8N force.
we can use vector addintion formula to fing the angle needed for the resulant of 4N and 6N to be 8N.
let x be the angle between the two forces 4N and 6N.
82=42+62-2*4*6*cos(x)
64-52=-24cos(x)
-1/2=cos(x)
therfore x=120degree